# Deep Linear Hawkes Processes

### v1

@@ 4 Related Works
Neural MTPPs Marked temporal point processes (MTPPs) are generative models that jointly model the time and type of continuous-time sequential events, typically characterized by mark-specific intensity functions (Daley & Vere-Jones, 2003). Early approaches, such as self-exciting Hawkes processes (Hawkes, 1971; Liniger, 2009), used simple parametric forms for the intensity. More recently, neural architectures such as RNNs (Du et al., 2016; Mei & Eisner, 2017), CNNs (Zhuzhel et al., 2023), and transformers (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) have been used to more flexibly model the conditional intensity. For intensity-free MTPPs, approaches include normalizing flows (Shchur et al., 2020a; Zagatti et al., 2024), neural processes (Bae et al., 2023), and diffusion models (Zeng et al., 2023; Zhang et al., 2024); however, the most common approach is to model intensities as it requires fewer modeling restrictions.
@@
Efficient MTPPs Due to their recurrent nature, RNN-based MTPP models incur $\mathcal{O}(N)$ complexity for sequences of length $N$ as events must be processed sequentially. Attention-based MTPP models can be applied in parallel across the sequence, but the computational work scales as $\mathcal{O}(N^{2})$. Türkmen et al. (2020) proposed modeling events as conditionally independent so long as they occurred within the same time bin of a specified size. This resulted in parallel computation within bins, but still scales overall as $\mathcal{O}(N)$. Shchur et al. (2020b) proposed an intensity-free TPP which uses triangular maps and the time-change theorem (Daley & Vere-Jones, 2003). This was extended by Zagatti et al. (2024) to handle marks, but in doing so, lost many of the benefits of the original model and scales linearly in the mark dimension—which can rapidly become untenable with $O(NK)$ work. To the best of our knowledge, our proposed model is the first that efficiently scales with sequence length and mark space, as well as the first to fully leverage SSMs and parallel scans.
@@
SSMs for Sequential Modeling SSMs have found recent success as alternatives to RNNs, CNNs, and transformers, enjoying reduced training cost and comparable modelling power (Gu et al., 2022b). A range of variants have been developed (Gu et al., 2021; Gupta et al., 2022; Gu et al., 2022a; Smith et al., 2022), and have been applied in language modeling (Gu & Dao, 2023), speech (Goel et al., 2022), and vision (Wang et al., 2023; Zhu et al., 2024). The linear recurrence allows for parallelism, as well as accessible long contexts which would be prohibitive for transformers due to their quadratic scaling. However, SSMs have not previously been applied to MTPPs, in part due to the irregular inter-event times and the input being a stochastic counting process.
Concurrent work by Gao et al. (2024) used Mamba (Gu & Dao, 2023), a recent deep SSM architecture, in an MTPP setting, which they call the Mamda Hawkes Process (MHP). The MHP uses a mamba SSM as the encoder in an encoder-decoder architecture, also leveraging the variable interval capabilities. Crucially, however, they use a separate parametric decoder for intermediate intensities (similar to, for instance, the THP). This misses the opportunity to “fully” leverage the SSM architecture, re-using the same variable interval evaluation to evaluate the the intensity.
5 Experiments
We now evaluate our deep linear Hawkes process model. Our core objectives in using SSMs for MTPP modeling were to define an architecture that is both (a) highly performant in its forecasting ability, and (b) able to leverage efficient parallel compute methods to accelerate inference. To this end, we first present a simple exploration of the ability of different models to represent parametric intensity functions. Then we present the main experiments in this paper, where we evaluate our model against a suite of common MTPP models on a range of datasets of different sizes. We conclude by testing the runtime of our model against a variety of baselines. We find that DLHP systematically outperforms baseline methods both in terms of log-likelihood on held-out test data and runtime across a range of sequence lengths. More results and details are included in Appendices C and D.
Metrics
[…]
Models
We compare our model (DLHP) with six of the most common MTPP models: two RNN-based models (RMTPP (Du et al., 2016), NHP (Mei & Eisner, 2017)), three transformer/attention-based models (THP (Zuo et al., 2020), SAHP (Zhang et al., 2020), AttNHP (Yang et al., 2022)), and one intensity-free model (IFTPP (Shchur et al., 2020a)). In all real-world experiments, extensive grid searches were conducted for hyperparameter tuning with configurations chosen based on validation log-likelihood, with the search spaces designed to roughly control parameter counts across models. Specifics for training, hyperparameters, and architectures are given in Appendix C.
@@ 5 Experiments
Table 1: Per event log-likelihood ($\uparrow$ is better) results on the held-out test set averaged over five seeds (standard deviations in parentheses); OOM indicates insufficient memory. We bold the best and underline the runner-up per dataset. We also report the mean rank of models across datasets as a summary metric ($\downarrow$ is better). DLHP is consistently the best or second best-performing model. Extended results and discussion are presented in Section D.1.
@@
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT
RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 6.3
@@
IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 2.9
DLHP (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 1.243 (0.083) -2.512 (0.369) 1.4
Libraries and Compute Environment
We implement our DLHP in the EasyTPP library (Xue et al., 2023) and use their implementations of the baseline models. We also use the five standard datasets that EasyTPP immediately supports (see Section C.2 for more details). We then further include three larger datasets to stress-test the MTPP models (see Section 5.2). Unless otherwise stated, all models were trained using a single NVIDIA A5000 GPU with 24GB of onboard memory.
5.1 Synthetic Experiments
We start by performing a simple investigation into the expressivity of the DLHP intensity function and the ability to capture background intensities. We train our model and baselines on 5,000 sequences over the time period $[0,7.5]$ drawn from an unmarked, inhomogeneous Poisson process with a square-wave intensity function, $\lambda_{t}:=\mathds{1}(t\in(1,2)\cup(3,4)\cup(5,6))$ (see Fig. 1). We plot the estimated intensity functions conditioned on no events occurring, i.e., $\mathcal{H}_{t}:=\emptyset\;\forall t$.
Intensity estimates are shown for NHP and DLHP specifically in Fig. 1 (and for all models in Fig. 6 in Section D.2). We can see that our model successfully captures the true, underlying background intensity process almost perfectly. This is largely attributed to the expressivity of the linear recurrences and non-linear depth of the model. Other models have various failure modes: struggling to capture the multi-modality (RMTPP, NHP, SAHP, and THP), not matching the square shape (previous four and IFTPP), or not being able to stop the pattern from repeating a fourth time (AttNHP).
We also examine the performance of our DLHP on randomly instantiated parametric Hawkes processes, finding that DLHP is able to successfully recover the ground truth intensity. Furthermore, in this setting DLHP achieves the best held-out log-likelihood scores and competitive next-event time prediction (see Section D.3 for more details). These simple experiments confirm that the DLHP is sufficiently expressive to be able to represent more complicated intensity functions while other methods break down.
5.2 Log-likelihood Results on Real-world Datasets
We empirically investigate the performance of our proposed model against baseline methods by comparing the held-out log-likelihood per event. We evaluate our model on eight real-world datasets. Five of which are taken directly from EasyTPP (Xue et al., 2023). We also include two MTPP datasets that have been widely used throughout the literature: Last.fm, which includes data on users’ music listening patterns from Celma Herrada et al. (2009), and MIMIC-II, a subset of de-identified patient hospital visits processed from (Saeed et al., 2002). Finally, we introduce a third dataset from the recently released, publicly available electronic health record (EHR) dataset EHRSHOT (Wornow et al., 2023). To construct the dataset, we first establish the most used Current Procedural Terminology (CPT-4) codes that identify medical services and procedures as events. The processed dataset comprises sequences of CPT-4 codes issued to individual patients during their care. This dataset has a maximum sequence length 10$\times$ longer than the longest in the EasyTPP datasets (and 100$\times$ that of MIMIC-II), providing a challenging testbed (in terms of scale) beyond existing datasets. Data statistics and other details including pre-processing are provided in Table 6 and Section C.3.
Figure 3: Per event log-likelihood on the held-out test data in Table 1 decomposed into time and mark components (i.e., $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{Time}}+\mathcal{L}_{\text{Mark}}$). Models are ordered by their average ranking. Model results are adjusted by subtracting the log-likelihood achieved by RMTPP for readability.
From the held-out log-likelihood results shown in Table 1, DLHP consistently achieves the best or the second-best log-likelihood across all datasets. Compared to the best baseline model per-dataset, DLHP produces a (geometric) mean likelihood ratio of 1.38 (corresponding to 38% higher likelihood on true events). We decompose this improvements in Fig. 3, where we see the improvements in log-likelihood are mainly driven by better modeling of time. Extended plots included in Section D.1. Given the clear improvement in temporal modeling, we posit that DLHPs are particularly well suited in applications that contain more complex patterns over time. All of these results for DLHP utilize input-dependent dynamics (see Section 3.4). This was found to reliably improve forecasting performance in ablation studies (see Section D.4).
We also report and discuss additional metrics in the Appendix. We report next-mark classification accuracy and RMSE of next-event arrival time, finding that DLHP matches or outperforms all baselines. We also report model calibration with respect to next event time and mark prediction. Calibration aims to grade the predictive uncertainty of the model (Bosser & Taieb, 2023), which is not captured by other metrics such as mark classification accuracy and time RMSE. On the whole, our model (as well as the baselines) tend to produce well-calibrated time and mark predictions across the datasets. We also include full tables for the likelihood decomposition in Table 7.
5.3 Speed Testing
A key motivation for DLHP is to leverage the properties of SSMs to improve scaling and accelerate inference. To test this, we measure the wallclock time for a full forward pass and log-likelihood evaluation on random input sequences with lengths ranging from eight events to over half a million events. The architectures and mark spaces are the same as in the StackOverflow experiments (see Tables 5(a) and 6). Our EasyTPP PyTorch DLHP is written in pure PyTorch, and hence is not as optimized as other methods (compared to, for instance, IFTPP, which uses a GPU-optimized implementation of the GRU). We therefore include the runtimes of a standalone JAX DLHP implementation, which allows for similar optimization through JIT compilation.
Figure 4: Median runtime, over 20 random seeds, for various models against increasing sequence lengths. We show runtimes for both conditioning on a sequence (Algorithm 1) and likelihood evaluation (Algorithm 3). We see that DLHP is faster across a wide range of sequence lengths. Variance of runtime is negligble. Crosses indicate where the THP runs out of memory or the IFTPP throws an error.
The THP scales well before reverting to superlinear scaling, and then runs out of memory. The IFTPP is very fast at shorter runtimes, but quickly reverts to linear scaling, due to its simple but highly optimized implementation and inherently sequential operation. Both DLHP implementations scale linearly at long sequence lengths, but have near-constant runtime at shorter sequences. At shorter sequence lengths, the more optimized JAX implementation is faster than the unoptimized pure PyTorch implementation. While this indicates that there are additional opportunities to accelerate the PyTorch implementation further (e.g., exploiting kernel fusion or writing a lower-level Triton implementation), these results still confirm our core hypothesis: that our DLHP can exploit parallel scans to scale to long sequences more effectively than other methods.

### v2

@@ 4 Related Work
Neural MTPPs: MTPPs are generative models that jointly model the time and type of continuous-time sequential events, typically characterized by mark-specific intensity functions (Daley and Vere-Jones, 2003). Early approaches used parametric intensity functions, such as self-exciting Hawkes processes (Hawkes, 1971; Liniger, 2009). More recently, neural models such as RNNs (Du et al., 2016; Mei and Eisner, 2017) and transformers (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022) were developed to enable flexible modeling of conditional intensities. Intensity-free MTPPs include normalizing flows (Shchur et al., 2020a; Zagatti et al., 2024), transformers (Draxler et al., 2025), neural processes (Bae et al., 2023), and diffusion models (Zeng et al., 2023; Zhang et al., 2024); however, modeling intensities is more common, requiring fewer modeling restrictions.
@@
Efficient MTPPs: Due to their recurrent nature, RNN-based MTPPs incur $\mathcal{O}(N)$ scaling for length-$N$ sequences as events are processed sequentially. Attention-based MTPP models can be applied in parallel across the sequence, but with $\mathcal{O}(N^{2})$ computational work. Türkmen et al. (2020) model events as conditionally independent if they occur within the same time bin of a specified size, resulting in parallel computation within bins, but still scaling as $\mathcal{O}(N)$ overall. Shchur et al. (2020b) proposed an intensity-free TPP using triangular maps and the time-change theorem (Daley and Vere-Jones, 2003). This was extended by Zagatti et al. (2024) to handle marks but losing benefits of the original model and scaling linearly in the mark dimension—which can rapidly become untenable as $\mathcal{O}(NK)$ work. Our S2P2 scales as $\mathcal{O}(\log N)$ and efficiently in marks; more discussion in Section˜B.5.
@@
SSMs for Sequential Modeling: SSMs have found recent success as alternatives to RNNs, CNNs, and transformers, enjoying reduced training cost and comparable modelling power (Gu et al., 2022a). A range of variants have been developed (Gu et al., 2021; Gupta et al., 2022; Gu et al., 2022b; Smith et al., 2022) and applied in language modeling (Gu and Dao, 2023), speech (Goel et al., 2022), and vision (Wang et al., 2023; Zhu et al., 2024). The linear recurrence enables parallelism, as well as accessible long contexts that are prohibitive for transformers due to quadratic scaling. However, SSMs have not previously been used as continuous-time models for MTPPs, in part due to the intensity functions having different left and right limits, and the input being a stochastic counting process.
SSMs for TPPs:Gao et al. (2024) propose using an SSM as a discrete sequence model, encoding an event sequence into a fixed set of static hidden states, then computing intensities with a separate parametric decoder (similar to transformer Hawkes processes (Zuo et al., 2020), but with the transformer replaced with Mamba (Gu and Dao, 2023)). This is fundamentally different to our continuous-time model, where instead of a separate parametric decoding head, we leverage the continuously evolving latent state at any time $t$ to compute predicted intensities at corresponding times. We empirically compare to their model MHP in Section˜6, finding that MHP (i) performs comparably to the THP, and (ii) is comprehensively outperformed by S2P2.
5 Synthetic Experiments
We first verify the ability of our S2P2 model to represent known intensity functions through a series of targeted experiments on synthetic data. Recent results from the deep SSM literature (Muca Cirone et al., 2024) show that sufficiently deep SSMs are able to approximate arbitrary continuous functions; their results apply to our S2P2 model, and hence we should be able to (verifiably) recover arbitrary and known intensity functions. To demonstrate this, we showcase recovering ground truth intensities in three different settings: the classical point processes of Hawkes and self-correcting processes; an inhomogeneous Poisson process with a discontinuous intensity function; and finally a marked process with long-range dependencies between marks. Our S2P2 recovers the correct intensity in all cases. We include full results for all baselines and reproducibility details for all the following examples and tasks in Appendix˜D, including intensity visualizations, where S2P2 works as intended, corrects recovers intensities, and matches or exceeds baseline methods performance.
Figure 3: Intensity estimates from trained models when conditioned on an empty sequence $\mathcal{H}_{t}=\emptyset$ for NHP (Mei and Eisner, 2017) and our S2P2. Dotted lines show the ground truth intensity for an inhomogeneous Poisson process. S2P2 accurately captures the background intensity.
Classical Point Processes: We first apply our model and baseline models to two classical point processes: a Hawkes process and a self-correcting process with intensity functions $\lambda_{t}=0.5+\sum_{i}0.5\exp(t_{i}-t)$ and $\lambda_{t}=\exp(t-N_{t})$, respectively. Examples of the recovered intensity functions are shown in Fig.˜4(a). We see that our S2P2 recovers the ground truth intensities nearly perfectly. We also explore applying our model and other baseline models to randomly generated synthetic multivariate Hawkes processes in Section˜D.3, finding that all methods perform comparably.
Inhomogeneous Poisson Process: Our S2P2 does not have a fixed parametric form for the intensity decoder (cf. THP or MHP) or limited recurrent dynamics (cf. NHP). This flexibility should allow the S2P2 to capture intensities where other methods fail. This is showcased in Fig.˜3. Models are trained on sequences drawn from an inhomogeneous Poisson process with a square wave for an intensity function (except for $t>7$). We observe that baseline models fail in predictable ways due to their expressivity limitations (please refer to Section˜D.2 for full results), whereas our model successfully captures the true background intensity process almost perfectly.
Long-Range Dependencies: Lastly, we study a task with known long-range dependencies, depicted in Fig.˜4(b). In this task, a “trigger” mark (green) is drawn from a homogeneous Poisson process ($\lambda$=0.1); each trigger is then followed by a “target” mark (orange) a predictably long time later. Here “long” refers to the fact that (a) there are often many trigger marks before a single target mark, and (b) the variance of the distribution over the trigger-target time is comparatively narrow compared to the mean. Gray marks are “distractors” drawn from a homogeneous Poison process ($\lambda$=1). We see that S2P2 successfully captures the long-range dependencies, whereas NHP struggles. In spite of the long delay between cause and effect within the data, S2P2 was able to recover 98% of the true data likelihood, whereas NHP only achieved 88%. This shows that the state-space point process successfully captures long-range dependencies and is flexible as an approximation family.
(a) S2P2 successfully recovers parametric intensities from self-correcting and Hawkes processes.
(b) Long-range dependencies between $\rightarrow||$ are accurately captured by S2P2, whereas NHP struggles to recover $\lambda_{t}$.
Figure 4: Results for synthetic experiments in Section˜5.
6 Real-World Experiments
We now present the main experiments in this paper, evaluating on real-world datasets. [^2] Our model is fully integrated into the EasyTPP (Xue et al., 2023) library [link]. Other code changes to reproduce our results can be found in our forked repository [link]. Model checkpoints are available on request. Further results and extensive experimental details are included in Appendices˜C and D.
Datasets:
We compare models on eight different datasets, including five datasets available from EasyTPP (Xue et al., 2023) (Amazon, Retweet, Taxi, Taobao and StackOverflow). We also add two commonly used datasets in the literature (Last.fm and MIMIC-II), as well as a new medical events dataset derived from the publicly available EHRSHOT dataset (Wornow et al., 2023). EHRSHOT is an order of magnitude larger than any dataset in EasyTPP in both the number of marks and the maximum sequence length (see summary statistics in Table˜7), representing a challenging application.
Metrics:
We performed a comprehensive evaluation using three common families of metrics: (i) total per-event log-likelihood (Mei and Eisner, 2017; Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022; Gao et al., 2024) (also broken down into mark- and time-specific likelihoods); (ii) predictive accuracy summary metrics of next-mark classification and next event time RMSE (Du et al., 2016; Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022; Gao et al., 2024); and (iii) mark and event time calibration metrics (Zhang et al., 2020; Bosser and Taieb, 2023) measuring the reliability of the implied uncertainties (see Section˜D.7). In order to characterize overall performance across multiple diverse metrics, we also compute a summary composite rank in Table 1(b), defined as the average rank obtained for a given model across all dataset-metric combinations across five random seeds (i.e., is an average of 240 individual evaluations and five differently initialized trained models).
Models:
We use the EasyTPP library (Xue et al., 2023) for baseline models and S2P2. We perform extensive hyperparameter searches for each individual pair of baseline model and dataset, then select the configurations that maximizes validation log-likelihood per model/dataset. The range of grid search and other details are elaborated in Section˜C.1; results are reported on a fully held-out test set. All models were trained on a single 24GB NVIDIA A5000 GPU.
@@ 6 Real-World Experiments
Table 2: Results for S2P2 and baselines across key metrics. We show the mean on the held-out test set evaluated across five random training seeds (std. dev. in parentheses). OOM indicates insufficient memory. We also report the mean rank of models across datasets as a summary metric (for which lower $\downarrow$ is better). S2P2 is consistently the best or second best-performing model on average on each dataset and metric. Extended results and discussion are shown in Section˜D.1.
(a) Per event total log-likelihood. Higher values indicate better performance.
Model Per Event Log-Likelihood, $\mathcal{L}_{\text{Total}}$ (nats) (Higher $\uparrow$ is better, x = best, x = second-best) Average
@@
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
RMTPP -2.136 (0.003) -7.098 (0.217) 0.346 (0.002) 1.003 (0.004) -2.480 (0.019) -1.780 (0.005) -0.472 (0.026) -8.081 (0.025) 7.1
SAHP -2.074 (0.029) -6.708 (0.029) 0.298 (0.057) 1.168 (0.028) -2.341 (0.058) -1.646 (0.083) -0.677 (0.072) -6.804 (0.126) 5.8
THP -2.096 (0.002) -6.659 (0.007) 0.372 (0.002) 0.790 (0.002) -2.338 (0.014) -1.712 (0.011) -0.577 (0.011) -7.208 (0.096) 6.1
IFTPP 0.496 (0.002) -10.344 (0.016) 0.453 (0.002) 1.318 (0.017) -2.233 (0.009) -0.492 (0.017) 0.317 (0.052) -6.596 (0.240) 3.0
MHP -2.091 (0.002) -6.564 (0.015) 0.370 (0.008) 0.636 (0.004) -2.346 (0.012) -1.676 (0.004) -0.351 (0.012) -7.206 (0.407) 5.9
@@
S2P2 (Ours) 0.781 (0.011) -6.365 (0.003) 0.522 (0.004) 1.304 (0.039) -2.163 (0.009) -0.557 (0.046) 0.919 (0.069) -2.512 (0.369) 1.4
(b) RMSE of the next event time prediction. Lower RMSE values indicate better performance.
Model Next Event Time RMSE (Lower $\downarrow$ is better, x = best, x = second-best) Average
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
RMTPP 0.338 (0.000) 16488 (70.5) 0.283 (0.001) 0.126 (0.000) 1.049 (0.003) 15.873 (0.000) 0.749 (0.010) 3425 (0.2) 5.0
SAHP 0.335 (0.001) 16102 (62.4) 0.290 (0.008) 0.126 (0.000) 1.031 (0.011) 15.757 (0.007) 1.142 (0.198) 3374 (9.4) 4.3
THP 0.332 (0.000) 16268 (18.7) 0.285 (0.001) 0.125 (0.000) 1.033 (0.005) 15.871 (0.000) 0.768 (0.005) 3414 (1.0) 4.3
IFTPP 0.327 (0.000) 16625 (0.2) 0.362 (0.178) 0.125 (0.000) 1.340 (0.724) 16.508 (0.555) 0.767 (0.029) 3616 (17.6) 5.0
MHP 0.329 (0.000) 16109 (36.9) 0.284 (0.003) 0.126 (0.000) 1.046 (0.030) 15.871 (0.000) 0.758 (0.065) 3418 (5.8) 4.0
NHP 0.339 (0.000) 15911 (4.0) 0.282 (0.001) 0.126 (0.000) 1.019 (0.001) 15.733 (0.008) 0.726 (0.001) 3330 (30.9) 2.4
AttNHP 2.656 (1.950) 16171 (284.2) 1.739 (0.422) 0.130 (0.000) 1.256 (0.030) 15.865 (0.017) 0.860 (0.022) OOM 6.6
S2P2 (Ours) 0.327 (0.000) 15987 (13.7) 0.281 (0.000) 0.126 (0.000) 1.014 (0.001) 15.720 (0.000) 0.894 (0.054) 3368 (14.4) 2.3
(c) Next mark classification accuracy (top-10 accuracy on EHRSHOT). Higher values indicate better performance.
Model Next Mark Classification Accuracy (%) (Higher $\uparrow$ is better, x = best, x = second-best) Average
Amazon Retweet Taxi Taobao StackOverflow Last.fm MIMIC-II EHRSHOT Rank ($\downarrow$)
RMTPP 30.8 (0.1) 53.4 (0.6) 91.4 (0.1) 60.9 (0.1) 45.6 (0.3) 52.5 (0.1) 92.3 (0.3) 36.5 (0.2) 6.9
SAHP 32.4 (1.0) 57.5 (2.2) 91.4 (0.7) 60.5 (0.2) 44.7 (2.0) 51.8 (0.7) 86.8 (0.9) 49.6 (3.1) 7.1
THP 34.6 (0.1) 60.2 (0.1) 91.4 (0.0) 60.0 (0.0) 46.6 (0.2) 53.3 (0.1) 90.9 (0.2) 49.7 (1.7) 5.6
IFTPP 35.9 (0.1) 50.4 (2.5) 91.8 (0.0) 61.0 (0.1) 45.6 (0.1) 56.4 (0.1) 93.4 (0.1) 62.7 (3.8) 4.3
MHP 35.1 (0.1) 60.0 (0.2) 91.4 (0.3) 60.7 (0.2) 46.5 (0.1) 54.3 (0.1) 93.2 (0.3) 54.8 (4.5) 5.0
NHP 39.4 (0.1) 61.4 (0.0) 92.9 (0.1) 61.5 (0.2) 47.1 (0.1) 56.5 (0.1) 94.3 (0.0) 76.5 (0.2) 1.8
AttNHP 38.9 (0.9) 60.7 (0.2) 92.6 (0.1) 61.3 (0.2) 48.2 (0.2) 55.8 (0.6) 92.9 (0.6) OOM 2.9
S2P2 (Ours) 40.7 (0.0) 61.3 (0.0) 93.1 (0.1) 61.1 (0.1) 47.5 (0.3) 55.8 (0.4) 96.0 (0.4) 79.5 (0.3) 1.9
Main Results: We report full numerical results in Table˜2, and a summary of rankings in Table 1(b). Table˜2(a) shows that S2P2 consistently achieves the best or the second-best held-out log-likelihood results across all datasets, beating other methods by over a whole rank and with a (geometric) mean likelihood ratio of 1.33 (corresponding to a 33% higher likelihood of true events). This is calculated by computing the mean log-likelihood ratio across all datasets and then exponentiating. We further investigate the log-likelihood improvement of S2P2 by separating the time- and mark-specific log-likelihood, finding that S2P2 is mainly driven by better temporal modeling and while achieving gains in both time and mark modeling over existing methods (see Tables˜8 and 6 in Section˜D.1 for full results). Additionally, we report prediction summary metrics, next event time prediction and next mark classification, in Table˜2(b) and Table˜2(c). We again see that S2P2 performs well, matching or exceeding the best-performing baseline, and far outpacing other baseline methods.
Additional Experiments: We defer several additional experiments and explorations to Appendices:
  1. 1.
    Calibration results: We also include full calibration results, both in raw numerical form and as reliability graphs (Bosser and Taieb, 2023). These show that S2P2 is well calibrated, placing highest amongst intensity-based methods, only being beaten by IFTPP (Shchur et al., 2020a).
  2. 2.
    Computational scaling: We define the theoretical time complexity of each baseline, where S2P2 is best with logarithmic complexity and linear work. We verify this scaling empirically.
  3. 3.
    Ablations: We also perform an ablation study on whether input dependent dynamics and the choice of which input to hold constant over the integration interval (introduced in Sections˜3.2 and 3.3) affect performance. We find input-dependent dynamics nearly always improve performance, and that performance is less sensitive to the specific form of the LLH’s input signal $\mathbf{u}_{t}$ in ZOH.
7 Discussion
Continuous-Time Hidden States:
The S2P2 and NHP (Mei and Eisner, 2017) architectures are closely related, in that both model continuous-time latent states. For the NHP, the intensity function is parameterized by a hidden state $\mathbf{h}(t)$ evolved via a continuous-time LSTM variant. Similarly, S2P2 continuously evolves a set of continuous-time latent states $\mathbf{h}^{1:L}(t)$, one for each layer $\{1,\ldots,L\}$. Then the hidden state $\mathbf{h}^{L}(t)$ of the top layer is decoded into intensities at any $t$. In contrast, models such as THP, MHP, IFTPP, only have latent states $\mathbf{h}(t_{j})$ at discrete event times $t_{j\in\{1:N\}}$. Consequently, some parametric shapes need to be defined to “interpolate” the hidden state values between events, thus sacrificing models’ flexibility due to these parametric functions.
Our results are grouped according to whether or not models have continuous-time hidden states. We empirically verify that continuously evolving hidden states enable more expressive latent dynamics, where the expressiveness in explaining the data is measured through log-likelihood; see Section˜D.1. S2P2 further improves scalability by incorporating SSMs that are naturally in continuous-time and by leveraging efficient parallelized scans. S2P2 also introduces a stronger inductive bias than most MTPP models (similar to NHP), helping to achieve superior next event predictions.
Limitation and Future Directions:
S2P2 forgoes the interpretability of the parameters and latent states of linear Hawkes processes (LHPs), meaning it is not appropriate when the interpretation of the underlying system is crucial. This tradeoff is common with neural MTPP models (e.g., NHP), because the limited expressivity of LHP may lead to poor predictive performance, as well as misleading interpretations due to underfitting. Therefore, an exciting opportunity for future research is exploiting the connection to recover LHP-like interpretability while retaining the enhanced predictive power of S2P2. For example, any recently developed techniques for interpreting deep SSMs can be extended to S2P2, such as Ali et al. (2025), while fully taking advantage of the continuously varying attention from continuous-time latent dynamics of S2P2 will provide richer insights compared to discrete-time attention maps (e.g., large language models).
Beyond this, further directions include leveraging theories and best practices from deep SSMs, such as the enhanced parameterizations presented by Merrill et al. (2024). S2P2 can also naturally be extended and accommodate non-categorical marks by tailoring its architecture to the structure of the mark space, such as marks having continuous values or containing richer information. A concrete example is spatio-temporal point processes that have applications in seismic and weather forecasting. Other valuable extensions include serving S2P2 as a pre-trained backbone for downstream applications, such as EHRSHOT clinical classification tasks. Additionally, censoring or adversarial conditions are commonly seen in practice (Boyd et al., 2023; Chakraborty et al., 2025). Extending the evaluation of S2P2 and MTPP models in general remains an open direction for real-world MTPP deployments.